OpenStudy (anonymous):

Solve 3x2 + 2x - 4 = 0 The 3x2 is 3x squared . Plz help :)

OpenStudy (anonymous):

Use quadratic formula: (-2 +/- SQRT(4 - 4*(-4)*3))/6 (-2 +/- SQRT(4+48))/6 (-2 +/- SQRT(52))/6

OpenStudy (anonymous):

\[3x^2 + 2x - 4 = 0\]\[(x+1/3)^2=13/9\]\[x=1/3(-1-\sqrt13)\]\[x=1/3(\sqrt13-1)\]